Air having a specific volume of 1.25 kg/m3 enters a gas turbine at a velocity of 105 m/s through an inlet area of 0.05 m2 . The air stream exits from the gas turbine at a velocity of 135 m/s and a specific volume of 0.67 kg/m3 . During the flow process, air loses 27 kJ/kg of heat and its specific enthalpy comes down by 145 kJ/kg. Determine, (i) the mass flow rate of air though the turbine, (ii) the turbine exit area, (iii) the power developed by the turbine


1
Expert's answer
2021-01-01T14:05:57-0500

i)


"\varDelta KE=\frac{v_2^2-v_1^2}{2}=3.6 kJ/kg"

ii)


"\frac {1}{\nu_1}A_1v_1=\frac {1}{\nu_2}A_2v_2"


"A_2=0.02 m^2"

iii)


"p=mW"


"m=\frac {1}{\nu_1}A_1v_1=4.2 kg"


"p=113.4 kJ"